# Claas Voelcker

### v1

@@ A.1 Main propositions
The first equality is due to the assumption that $V$ is the true value function. Using the fact that $\hat{x}^{(1)}_{i}$ and $x^{(1)}_{i}$ are sampled from the same distributions, but are conditionally independent, and using $V^{\prime}=V$, we can state
$\mathbb{E}\left[V(\hat{x}^{(1)}_{i})V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\right]=\mathbb{E}\left[\mathbb{E}\left[V(\hat{x}^{(1)}_{i})\middle|x^{(0)}_{i}\right]\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)|x^{(0)}_{i}\right]\right]=\mathbb{E}\left[V(\hat{x}^{(1)}_{i})^{2}\right].$
Using this fact, we obtain:
$\mathbb{E}\left[\left[V(\hat{x}^{(1)}_{i})\right]^{2}-2V(\hat{x}^{(1)}_{i})V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)+V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)^{2}\right]$
$=\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)^{2}-V(\hat{x}^{(1)}_{i})V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\right]$
$=\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)^{2}\right]-\mathbb{E}\left[V(\hat{x}^{(1)}_{i})V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\right]$
$=\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)^{2}\right]-\mathbb{E}\left[\mathbb{E}\left[V\left(\hat{x}^{(1)}_{i}\right)\middle|x^{(0)}_{i}\right]\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\middle|x^{(0)}_{i}\right]\right]$
$=\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)^{2}\right]-\mathbb{E}\left[V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\right]^{2}.$
Therefore, at $\hat{p}=p$
$\mathbb{E}\left[\hat{\mathcal{L}}_{\text{MuZero}}(p,V)\right]=\text{IterVAML}+\mathbb{E}\left[\text{Var}\left(V_{\text{est}}\left(x^{(1)}_{i},x^{(2)}_{i}\right)\middle|x^{(0)}_{i}\right)\right].$
The first term is minimized by the ground-truth model, but not the second term, which means the ground-truth model does not provide a value-equivalent model under the MuZero loss.
$\Box$
A.2 Propositions from Bertsekas and Shreve [1978]
For convenience, we quote some results from Bertsekas and Shreve [1978]. These are used in the proof of Lemma 1.
Proposition 3 (Proposition 7.30 of Bertsekas and Shreve [1978])
Let $\mathcal{X}$ and $\mathcal{Y}$ be separable metrizable spaces and let $q(\mathrm{D}y|x$ be a continuous stochastic kernel on $\mathcal{Y}$ given $\mathcal{X}$. If $f\in\mathcal{C}(\mathcal{X}\mathcal{Y})$, the the function $\lambda:\mathcal{X}\rightarrow\mathbb{R}$ defined by
$\lambda(x)=\int f(x,y)q(\mathrm{d}y|x)$
is continuous.
Proposition 4 (Proposition 7.33 of Bertsekas and Shreve [1978])
[…]
$f^{*}(x)=\min_{y\in\mathcal{D}_{\mathcal{X}}}f(x,y).$
[…]
@@ A.2 Propositions from Bertsekas and Shreve [1978]
In our proof, we construct $f^{*}$ as the minimum of an IterVAML style loss and equate $\phi$ with the function we call $f$ in our proof. The change in notation is chosen to reflect the modern notation in MBRL, in the textbook, older notation is used.
A.3 Bias due to the stabilizing loss
@@
As highlighted in section 3, the addition of a stabilizing loss is necessary to achieve good performance with any of the loss functions. With deterministic models, the combination ^$\hat{\mathcal{L}}_{\text{IterVAML}}+\hat{\mathcal{L}}^{n}_{\text{latent}}$ is stable, but the conditions for recovering the optimal model are not met anymore. This is due to the fact that $\operatorname*{arg\,min}_{\hat{f}}\mathbb{E}\left[\hat{\mathcal{L}}(\hat{(}f),\phi;\mathcal{D})\right]=\mathbb{E}\left[\phi(x^{\prime}\right]$, but in general $\mathbb{E}\left[V(\phi(x^{\prime}))\right]\neq V\left(\mathbb{E}\left[\phi(x^{\prime})\right]\right)$. While another stabilization technique could be found that does not have this problem, we leave this for future work.
Appendix B Implementation Details
B.1 Environments
@@
The experiments in the main body are conducted with the uniform noise formulation, the evaluation over the Gaussian mixture noise is presented in Appendix C.
B.2 Algorithm and pseudocode
The full code is provided in the supplementary material and will be publicly released on Github after the review period is over. A pseudo-code for the IterVAML implementation of $\lambda$-AC is provided for reference (algorithm 1). Instead of taking the simple $n$-step rollout presented in the algorithm, we use an average over all bootstrap targets up to $n$ for the critic estimate, but used a single $n$-step bootstrap target for the actor update. We found that this slightly increases the stability of the TD update. A TD-$\lambda$ procedure as used by Hafner et al. [2020] did not lead to significant performance changes.
@@ B.2 Algorithm and pseudocode
B.3 Model architecture and hyperparameter choices
@@
Similar to previous work in MBRL [Janner et al., 2019] we slowly increased the length of model rollouts for actor and critic training following a linear schedule over the first 25.000 steps. All hyperparameters are detailed in Table 1. Where relevant, variables refer to those used in algorithm 1 for clarity.
All neural networks are implemented as two or three layer MLPs, adapted from Hansen et al. [2022]. We found that minor architecture variations, such as adapting the recurrent architecture from Amos et al. [2021] did not have a large impact on the performance of the algorithm in the evaluated environments. Similarly, adding regularization such as Layer Norm or Batch Norm to the model as in Paster et al. [2021] or Ghugare et al. [2023] did not change the outcome of the experiments noticeably. We did keep the LayerNorm in the critic and actor input, as it had some stabilizing effect for the value-aware losses. The full impact of regularization in end-to-end MBRL architectures is an open question that deserves future study.
@@
Appendix C Additional experiments and ablations
C.1 Runtime estimate
Computationally efficiency was estimated on a dedicated machine with a RTX 2090 and rounded. Exact runtimes can vary depending on hardware setup. Overall we find that the $\lambda$-AC framework obtains similar runtime efficiency as related approaches which also differentiate the model for policy gradients [Ghugare et al., 2023] and is slightly slower than just using the model for MPC [Hansen et al., 2022]. Using the IterVAML loss requires another network pass through the value function compared to MuZero, therefore it is slightly slower in our current setup. However, these differences can most likely be improved with a more efficient implementation.
All model-based algorithms have a slowdown of roughly the model depth over SAC (×$\times 4-6$). This is expected, since the model is rolled out sequentially and the complexity of the computational graph grows linearly with the model rollout depth. Since SAC requires large networks and high UTD to obtain similar performance to the model-based approaches, MBRL is still runtime competitive to model-free RL (compare Ghugare et al. [2023]).
@@ C.1 Runtime estimate
$\lambda$-AC (IterVAML) $\lambda$-AC (MuZero) $\lambda$-AC (BYOL ablation) SAC
@@
C.2 Gradient propagation in IterVAML
Figure 8: Comparison of using the TD gradient for model learning as well in IterVAML $\lambda$-AC.
@@ C.2 Gradient propagation in IterVAML
C.3 Further experiments in noisy environments
@@
C.4 Removing latent formulation and stabilizing loss
@@
Finally, we present the results of removing both the latent and the auxiliary stabilization in all $\lambda$-AC algorithms in Figure 10. IterVAML diverges catastrophically, with value function error causing numerical issues, which is why the resulting graph is cut off: no run in 8 seeds was able to progress beyond 100.000 steps. This is similar to results described by Voelcker et al. [2022]. MuZero on the other hand profits from not using a stabilizing BYOL loss when operating in observation space. This substantiates claims made by Tang et al. [2022], which conjecture that latent self-prediction losses such as BYOL can capture good representation of MDPs for reinforcement learning. Therefore achieving a lower BYOl loss might be more a better auxiliary task than achieving a small MSE in observation space.

### v3

@@ A.3 Propositions from Bertsekas & Shreve (1978)
In our proof, we construct $f^{*}$ as the minimum of an IterVAML style loss and equate $\phi$ with the function we call $f$ in our proof. The change in notation is chosen to reflect the modern notation in MBRL – in the textbook, older notation is used.
A.4 Bias due to the stabilizing loss
@@
As highlighted in Subsection 3.1, the addition of a stabilizing loss is necessary to achieve good performance with any of the loss functions. With deterministic models, the combination ^$\hat{\mathcal{L}}_{\text{IterVAML}}+\hat{\mathcal{L}}^{n}_{\text{latent}}$ is stable, but the conditions for recovering the optimal model are not met anymore. This is due to the fact that $\operatorname*{arg\,min}_{\hat{f}}\mathbb{E}\left[\hat{\mathcal{L}_{\text{latent}}}(\hat{f},\phi;\mathcal{D})\right]=\mathbb{E}\left[\phi(x^{\prime}\right]$, but in general $\mathbb{E}\left[V(\phi(x^{\prime}))\right]\neq V\left(\mathbb{E}\left[\phi(x^{\prime})\right]\right)$. While another stabilization technique could be found that does not have this problem, we leave this for future work.
A.5 Multi-step IterVAML
In Subsection 2.1 the multi-step extension of IterVAML is introduced. As MVE and SVG require multi-step rollouts to be effective, it became apparent that simply forcing the one-step prediction of the value function to be correct is insufficient to obtain good performance. We therefore extended the loss into a multi-step variant.
Using linear algebra notation for simplicity, the single-step IterVAML loss enforces
$\min\left|\langle V,P(x,a)-\hat{P}(x,a)\rangle\right|^{2}$
The $n$-step variant then seeks to enforce a minimum between $n$ applications of the respective transition operators
$\min\left|\langle V,P^{n}(x,a)-\hat{P}^{n}(x,a)\rangle\right|^{2}$
The sample-based variant is a proper regression target, as $V(x^{(j)}$ is an unbiased sample from $P^{n}(x,a)$. It is easy to show following the same techniques as used in the proofs of propositions 1 and 2 that the sample-based version indeed minimizes the IterVAML loss in expectation.
Finally, we simply sum over intermediate $n$-step versions which results in the network being forced to learn a compromise between single-step and multi-step fidelity to the value function.
Appendix B $\lambda$-Reinforcement Learning Algorithms
Model loss Value est. Policy est. Actor policy DAML Latent
$\lambda$-IterVAML BYOL MVE SVG direct
$\lambda$-MuZero BYOL MuZero SVG direct
MuZero (Schrittwieser et al., 2020) - MuZero - MCTS
Eff.-MuZero (Ye et al., 2021) - MuZero - MCTS
ALM (Ghugare et al., 2023) ALM-ELBO m.-free ALM-SVG direct
TD-MPC (Hansen et al., 2022) BYOL m.-free DDPG MPC
MBPO (Janner et al., 2019) MLE SAC SAC direct - -
Dreamer (Hafner et al., 2020) ELBO MVE SVG direct -
IterVAML (Farahmand, 2018) - Dyna - direct -
VaGraM (Voelcker et al., 2022) weigthed MLE SAC SAC direct ✓(?) -
Table 1: An overview of different model-based RL algorithms and how they fit into the $\lambda$ family. The first two are the empirical algorithms tested in this work. The next section contains work that falls well into the $\lambda$ family as described in this paper. The similarities between these highlight that further algorithms are easily constructed, i.e. ALM (Ghugare et al., 2023) combined with MPC. The final section contains both popular algorithms and closely related work which inform the classification, but are not part of it since they are either not latent or not decision-aware.
We introduce the idea of the $\lambda$ designation in Section 3. The characterization of the family is fairly broad, and it contains more algorithms than those directly discussed in this paper. An overview of related algorithms is presented in Table 1. However, we found it useful to establish that many recently proposed algorithms are closely related and contain components that can be combined freely with one another. While the community treats, i.e. MuZero, as a fixed, well-defined algorithm, it might be more useful to treat it as a certain set of implementation and design decisions under an umbrella of related algorithms. Given the findings of this paper for example, it is feasible to evaluate a MuZero alternative which simply replaces the value function learning algorithm with MVE, which should be more robust in stochastic environments.
A full benchmarking and comparison effort is out of scope for this work. However, we believe that a more integrative and holistic view over the many related algorithms in this family is useful for the community, which is why we present it here.
Appendix C Implementation Details
C.1 Environments
@@
The experiments in the main body are conducted with the uniform noise formulation, the evaluation over the Gaussian mixture noise is presented in Appendix D.
C.2 Algorithm and pseudocode
The full code is provided in the supplementary material and will be publicly released on Github after the review period is over. A pseudo-code for the IterVAML implementation of $\lambda$ is provided for reference (algorithm 1). Instead of taking the simple $n$-step rollout presented in the algorithm, we use an average over all bootstrap targets up to $n$ for the critic estimate, but used a single $n$-step bootstrap target for the actor update. We found that this slightly increases the stability of the TD update. A TD-$\lambda$ procedure as used by Hafner et al. (2020) did not lead to significant performance changes.
@@ C.2 Algorithm and pseudocode
C.3 Model architecture and hyperparameter choices
@@
Similar to previous work in MBRL (Janner et al., 2019) we slowly increased the length of model rollouts for actor and critic training following a linear schedule over the first 25.000 steps. All hyperparameters are detailed in Table 2. Where relevant, variables refer to those used in algorithm 1 for clarity.
All neural networks are implemented as two or three layer MLPs, adapted from Hansen et al. (2022). We found that minor architecture variations, such as adapting the recurrent architecture from Amos et al. (2021) did not have a large impact on the performance of the algorithm in the evaluated environments. Similarly, adding regularization such as Layer Norm or Batch Norm to the model as in Paster et al. (2021) or Ghugare et al. (2023) did not change the outcome of the experiments noticeably. We did keep the LayerNorm in the critic and actor input, as it had some stabilizing effect for the value-aware losses. The full impact of regularization in end-to-end MBRL architectures is an open question that deserves future study.
@@
Appendix D Additional experiments and ablations
D.1 Runtime estimate
Computationally efficiency was estimated on a dedicated machine with a RTX 2090 and rounded to the closest 5 minute mark. Exact runtimes can vary depending on hardware setup. Overall we find that the $\lambda$ algorithms obtains similar runtime efficiency as related approaches which also differentiate the model for policy gradients (Ghugare et al., 2023) and is slightly slower than just using the model for MPC (Hansen et al., 2022). Using the IterVAML loss requires another network pass through the value function compared to MuZero, therefore it is slightly slower in our current setup. However, these differences can most likely be improved with a more efficient implementation with more clever reuse of intermediate computation.
All model-based algorithms have a slowdown of roughly the model depth over SAC (×$\times 4-6$). This is expected, since the model is rolled out sequentially and the complexity of the computational graph grows linearly with the model rollout depth. Since SAC requires large networks and high UTD to obtain similar performance to the model-based approaches, MBRL is still runtime competitive to model-free RL (compare Ghugare et al. (2023)).
@@ D.1 Runtime estimate
$\lambda$-IterVAML $\lambda$-MuZero BYOL ablation SAC
@@
D.2 Gradient propagation in IterVAML
Figure 8: Comparison of using the TD gradient for model learning as well in $\lambda$-IterVAML.
@@ D.2 Gradient propagation in IterVAML
D.3 Further experiments in noisy environments
@@
D.3.1 Tabular benchmark
Figure 10: Visualization of the value approximation error using tabular models with differing rank. Shaded area represents three standard errors of the mean.
As discussed, no widely used large-scale benchmark for stochastic MDPs exists. Therefore, we created a small scale benchmark based on the Garnet MDP framework (Archibald et al., 1995) as follows:
We constructed a deterministic Markov Reward Process with a finite state space consisting of states ($|\mathcal{S}|=n$). Its transition matrix is a permutation of the identity matrix. The reward was randomly sampled from a standard normal distribution and an additional reward of 10 was assigned to the first state to further increase the variance of the reward over all states, as this is the critical scenario for MuZero’s bias.
To introduce stochasticity, we allowed for $m-1$ additional successor states for each state $\mathcal{S}_{x}\subset\mathcal{S}$. We also introduced an additional parameter $\rho$ to interpolate between the deterministic Markov Reward Process and the stochastic one. The first successor state taken from the identity matrix was assigned the transition probability $\rho$ and all additional successor states were assigned probability $(1-\rho)/(m-1)$. This allowed us to control the stochasticity of the problem simply by varying $\rho$.
Since finite-state, deterministic models are non differentiable, we instead modeled the impact of the model class bias by constraining the capacity of the model. The model class is constructed from low rank factorized matrices, represented by a random representation matrix $\Phi$ and a learnable weight matrix $\Psi$ of dimensions $(n,k)$, so that $\hat{P}=\sigma(\Phi^{T}\Psi)$. The probability constraint are ensured via the row-wise softmax operation $\sigma$. By reducing k, the capacity of the model is constrained so that a correct model cannot be represented.
We did not constrain the representation of the value function and simply learned tabular value functions using fitted value iteration with a model-based target estimate and with the MuZero value function learning scheme. We used the original formulation of each loss as accurately as possible, with some minor adaptation for MuZero. This was necessary as the end-to-end gradient through the value function is impossible to compute with discrete samples. Instead, we used the difference between the expected value over the next state and the bootstrapped target value.
We conducted our experiment over 16 randomly generated garnets, rewards and representation matrices with $n=50$ and $m=10$. The data was sampled from a uniform distribution and next states according to the transition matrix. The reward function was not learned to simplify the setup. We sampled sufficient data points $(100,000)$ to ensure that an unconstrained MLE model can be learned to high accuracy. The model and value function were learned with full batch gradient descent using Adam. All experiments were simple value estimation for a fixed reward process with constrained model classes; we did not address policy improvement. We varied the constraint k between 10 and 1 to assess the impact of reduced model capacity on the algorithms, with varying $\rho\in[0.5,0.75,1.0]$. We also added experiments for $k=50$ to assess an unconstrained model, although in most cases, performance was already close to optimal with $k=10$.
The results of these experiments are presented in Figure 10, raw data in Table 5.
We found that for all values of stochasticity, the IterVAML model performs best for value function learning. Curiously, in all environments it is matched closely or outperformed slightly by the MLE solution with extremely low values of $k$, which suggests that the model does not have sufficient flexibility to model any value estimate correctly. In these cases, the MLE can provide a more stable learning target, even though none of the evaluated algorithms will be useful under extreme constraints.
Curiously, even in the deterministic case, MuZero does not perform well unless the model capacity is large enough. We investigated this further and found that due to our constraint setup, the model is unable to fully capture deterministic dynamics in the constrained case and will instead remain stochastic. Therefore the bias of the MuZero loss still impacts the system, as the variance over next state values induced by the model is high. We also noticed that the gradient descent based optimization of value function and model in MuZero has a tendency to converge to suboptimal local minima, which might explain some of the extreme outliers.
When replacing the value learning scheme with model-based TD learning, we recover almost identical performance between IterVAML and MuZero, which further highlights the results presented in this paper and by Grimm et al. (2021) that both losses optimize a similar target.
Overall, these experiments provide further evidence that the MuZero value learning scheme is flawed in stochastic environments, and with stochastic models.
$\rho=0.5$ IterVAML MuZero MLE
1 0.9406 +/- 0.0955 3.6476 +/- 0.11205 0.8417 +/- 0.0532
2 0.9149 +/- 0.02915 3.6486 +/- 0.115875 0.8671 +/- 0.035975
3 0.8949 +/- 0.07915 3.5668 +/- 0.1151 0.7687 +/- 0.031925
4 0.5728 +/- 0.0723 3.5281 +/- 0.123425 0.6941 +/- 0.05015
5 0.29 +/- 0.023825 3.3755 +/- 0.136025 0.5256 +/- 0.0268
6 0.2325 +/- 0.01965 3.1761 +/- 0.116925 0.5618 +/- 0.046075
7 0.174 +/- 0.02195 3.1889 +/- 0.160375 0.4792 +/- 0.047225
8 0.1391 +/- 0.0179 2.9965 +/- 0.12555 0.3936 +/- 0.038925
9 0.1032 +/- 0.0139 2.8649 +/- 0.11805 0.3623 +/- 0.029225
10 0.0577 +/- 0.005925 2.8381 +/- 0.151475 0.273 +/- 0.022625
50 0.0017 +/- 0.000225 2.5494 +/- 0.115075 0.0017 +/- 0.0002
$\rho=0.75$ IterVAML MuZero MLE
1 2.8279 +/- 0.283675 5.1395 +/- 0.257475 2.6313 +/- 0.210275
2 2.7452 +/- 0.1289 5.1149 +/- 0.27055 2.76 +/- 0.1613
3 1.8714 +/- 0.195725 4.8157 +/- 0.27655 2.4053 +/- 0.153125
4 0.7264 +/- 0.043275 4.5975 +/- 0.2393 1.6601 +/- 0.1318
5 0.3794 +/- 0.026425 4.3195 +/- 0.2547 1.1971 +/- 0.042725
6 0.2066 +/- 0.016675 3.4099 +/- 0.24005 0.9698 +/- 0.0657
7 0.1213 +/- 0.013075 3.1378 +/- 0.231275 0.5805 +/- 0.05625
8 0.1086 +/- 0.006675 3.4061 +/- 0.167275 0.6449 +/- 0.080075
9 0.0799 +/- 0.009 3.0232 +/- 0.23425 0.4361 +/- 0.042275
10 0.0456 +/- 0.0036 3.2613 +/- 0.29505 0.2999 +/- 0.028375
50 0.0027 +/- 0.00025 2.5132 +/- 0.194225 0.0026 +/- 0.000275
$\rho=1.$ IterVAML MuZero MLE
1 14.0457 +/- 1.50455 16.5522 +/- 1.79805 14.2642 +/- 1.527775
2 13.0414 +/- 1.100675 16.9774 +/- 1.636875 13.7411 +/- 1.30145
3 6.1846 +/- 0.688375 11.9329 +/- 1.19215 9.5593 +/- 1.193025
4 1.1092 +/- 0.219475 3.335 +/- 0.7467 2.248 +/- 0.355225
5 0.0977 +/- 0.026175 0.3616 +/- 0.107525 0.2503 +/- 0.08265
6 0.01 +/- 0.0031 36.276 +/- 23.983825 0.0042 +/- 0.001
7 0.0065 +/- 0.0028 0.0134 +/- 0.001125 0.0052 +/- 0.001325
8 0.0083 +/- 0.0021 0.0128 +/- 0.001275 0.0045 +/- 0.001475
9 0.0032 +/- 0.0004 0.0117 +/- 0.001075 0.0021 +/- 0.000425
10 0.0054 +/- 0.001675 0.0092 +/- 0.00105 0.0035 +/- 0.000725
50 0.0023 +/- 0.00035 0.0048 +/- 0.0013 0.0044 +/- 0.000775
Table 5: Garnet experiment results. Each cell depicts mean and standard deviation of value estimation over 16 randomly generated garnet MDPs for each algorithm. Minimum value highlighted (when ambiguous due to confidence interval, multiple values are highlighted). For the case of $\rho=1$, all losses achieve strong performance with $k\geq 7$, therefore no highlighting was necessary.
D.4 Removing latent formulation and stabilizing loss
@@
Finally, we present the results of removing both the latent and the auxiliary stabilization in all $\lambda$ algorithms in Figure 11. IterVAML diverges catastrophically, with value function error causing numerical issues, which is why the resulting graph is cut off: no run in 8 seeds was able to progress beyond 100.000 steps. This is similar to results described by Voelcker et al. (2022). MuZero on the other hand profits from not using a stabilizing BYOL loss when operating in observation space. This substantiates claims made by Tang et al. (2022), which conjecture that latent self-prediction losses such as BYOL can capture good representation of MDPs for reinforcement learning but reconstruction accuracy might not be a useful goal.